OpenStudy (anonymous):

@yummydum A machinist creates a washer by drilling a hole through the center of a circular piece of metal. If the piece of metal has a radius of x + 10 and the hole has a radius of x + 6, what is the area of the washer?

OpenStudy (anonymous):

fins the area of the metal piece then of the hole then subtract the area of the hole from the one of the metal piece

OpenStudy (anonymous):

the answer?

OpenStudy (anonymous):

lol....arent you going to work this one? :P

OpenStudy (anonymous):

not really lol :)

OpenStudy (anonymous):

\[A _{circle}=\pi r^{2}\]

OpenStudy (anonymous):

\[[\pi (x+10)^{2}]-[\pi(x+6)^{2}]\]

OpenStudy (anonymous):

try working it :)

OpenStudy (anonymous):

once you solve that thing i just wrote BOOM thats your answer :)

OpenStudy (anonymous):

But how do you solve it?????

OpenStudy (anonymous):

\[[\pi \times (x+10)(x+10)] - [\pi \times (x+6)(x+6)]\]\[[\pi \times (x^{2}+20x+100)] - [\pi \times (x^{2}+12x+36)]\]

OpenStudy (anonymous):